OpenStudy (anonymous):

Are the lines 2y-x=6 and 6x-3y-33=0 parallel , perpendicular, or neither?

OpenStudy (anonymous):

to see if the are parallel or perpendicular, you need to know the gradient of both lines. the best way to do this is to get it into the form y=mx+b where m is the gradient. So rearrange both formulas to get y by itself

OpenStudy (anonymous):

Okay , thank you.

OpenStudy (anonymous):

if they are parallel, both m's should be equal, if they are perpendicular m1*m2=-1 (m1 is the gradient of the 1st eqn and m2 is the second equation)